# Rotatable Array-Aided Hybrid Beamforming for Integrated Sensing and Communication

### v1

@@ III-F Overall Algorithm
The total complexity of Algorithm 2 stems from the solving the variables $\{\mathbf{u},\mathbf{W},\mathbf{F},\mathbf{\Lambda},\mathbf{\mu},\bm{\xi}^{c},\bm{\xi}_{s}\}$. The computational complexity for calculating receive beamformer $\mathbf{u}$ is $\mathcal{O}(N_{r}^{3})$ due to the matrix inversion in (24). Similarly, The computational complexity for optimizing transmit digital beamformer $\mathbf{W}$ and transmit analog beamformer $\mathbf{F}$ is both $\mathcal{O}(N_{t}^{3})$. To update the auxiliary variables $\mathbf{\mu},\bm{\xi}^{c}$ and $\bm{\xi}_{s}$, the complexities are $\mathcal{O}(KN_{t})$, $\mathcal{O}(KN_{t}+K^{2})$ and $\mathcal{O}(N_{t}K^{2})$, respectively. The complexity for updating $\mathbf{\Lambda}$ mainly stems from the derivative calculations (51) for the $L$ elements in $\mathbf{\Lambda}$. Thus, the computational complexity of $\mathbf{\Lambda}$ is $\mathbf{O}(I_{GA}LN_{t}^{2}K^{2})$, where $I_{GA}$ is the number of iteration for gradient ascent. Considering the domainant computational steps, the total computational complexity of Algorithm 2 is $\mathcal{O}(I_{AO}I_{GA}LN_{t}^{2}K^{2})$, where $I_{AO}$ denote the number of iterations of AO algorithm.
@@ IV Numerical results
In this section, we will elaborate on the simulation stetup and discussion of the numerical results to evaluate the performance of the proposed algorithm.
@@ IV-A Simulation Setup
In the simulation, we set the BS as the origin of the global coordinate system. The center coordinates of the transmitting uniform planar array (UPA) and the receiving UPA are set to $0,0,-d_{0}]^{T}$ and $0,0,d_{0}]^{T}$, respectively. We assume the BS is equipped with $N_{t}$ trainsmit antennas, $N_{r}$ receive antennas, and $N_{RF}$ RF chains. Each RF chain is connected to $M$ antennas via analog phase shifters. The antenna’s positions are set according to the transmit and receive UPAs with the largest achievable apertures $D\times D$. For each UPA, the plate size is $D\times D$ with $D=10\lambda$. The BS operates at a carrier frequency of 30GHz. It serves $K$ = 4 users, sensing $s=1$ target, and there are C=C=3 clutters in the scenario. The users and clutters are randomly distributed in front of the BS, i.e., $\phi_{d}\sim\mathcal{U}(-\frac{\pi}{2},\frac{\pi}{2})$, and $\theta_{d}\sim\mathcal{U}(0,\pi)$ for d $\in\{1,\dots,K+C\}$. The sensing target is assumed to be locate at $(\frac{\pi}{3},\frac{\pi}{3})$. The reflection coefficients of both the target and the clutter , $\zeta_{k}$ and $\zeta_{s}$ , are set to 2 dB. In addition, the power budget at the BS is set as $P_{max}$ = 0 dBm. The average noise power as $\sigma^{2}_{k}=\sigma^{2}_{s}=$ -80 dBm. The penalty factor $\eta=$ 10. The preliminary step size $\kappa=100$ and the minimum setp size $\kappa_{min}$ = 0.01. It is worth noting that, unless otherwise specified, the rotation variable $\mathbf{\Lambda}$ we optimized includes all rotation parameters of the Tx and Rx, i.e., $\mathbf{\Lambda}=[\alpha^{\mathrm{t}},\beta^{\mathrm{t}},\gamma^{\mathrm{t}},\alpha^{\mathrm{r}},\beta^{\mathrm{r}},\gamma^{\mathrm{r}}]$. We perform 200 Monte Carlo simulations for the following results. The simulation parameters for the proposed system are listed in Table. I
@@
To verify the effectiveness of the proposed scheme with Algorithm 2, we introduce five baselines to compare the performance improvedment of our algorithm, which are: 1) Tx and Rx have fixed rotation angles, with RF chains adopting a sub-connected structure (FPA, Sub-connected), 2) Tx and Rx have fixed rotation angles, with RF chains adopting a fully-connected structure (FPA, Fully-connected), 3) Tx and Rx have fixed rotation angles, using a fully-digital beamformer (FPA, Fully-digital), 4) Tx and Rx can execute 3D rotations, with RF chains adopting a fully-connected structure (RMA, Fully-connected),
@@ IV-B Convergence Evaluation
First, we evaluate the convergence of the proposed AO algorithm under different numbers of RF chains in Fig. 2. As can be seen, the algorithm converges rapidly in all scenarios, and the objective value tends to stabilize after 10 iterations. Moreover, the WSR increases as the number of RF chains increases. This is because the performance of the analog beamformer $\mathbf{F}$ is constrained by the number of RF chains. Specifically, when the number of RF chains increases, the analog beamformer can achieve a higher DoFs to further enhance the overall system performance.
@@ IV-C Beamfocusing of Algorithm 2
To verify the effect of Algorithm 2 on improving sensing performance, we consider the optimized beampattern by Algorithm 2 and compares it with the optimized beampattern by FPA scheme. This is to intuitively demonstrate the advantages of rotatable array in the sensing function of ISAC system proposed in this paper. The beampattern at the angle ($\phi,\theta$) is computed as
@@
We set $\omega_{s}=1$ to eliminate the interference from CUs. The number of transmit antennas and receive antennas are both $N_{t}=N_{r}=16$, the number of RF chains is $N_{RF}=4$, and the target is assumed to be located at ($60^{\circ},60^{\circ}$). Figs.3(a) shows the optimized beampattern under the FPA scheme, from which it can be seen that the beam fails to focus well on the target position. In Figs.3(b), the side length of the UPA is set to $D=5\lambda$, and it is observed that under the RA-aided scheme, the beam can focus precisely on the desired target position with extremely low interference to other positions. In the far-field scenario, as the side length of the UPA increases as shown in Fig.3(c), the width of the beam main lobe narrows gradually, however, due to the increase in the spacing between antennas in the array, the array periodicity is enhanced, the number of grating lobes increases, and thus more interference is generated.
@@
Figure 3: Beampattern under different configurations
@@ IV-D Simulation Results
Fig. 4 compares the relationship between the WSR and the maximum transmit signal power $P_{max}$ among different schemes. As shown in the figure, with the increase of transmit signal power, the WSR values of both the RA and FPA schemes keep rising, which is because a higher transmit power can enhance the signal strength. Under the condition that both the transmit power and precoding structure are the same, the performance of the RA scheme is consistently superior to that of the FPA scheme, which is attributed to the higher flexibility of the RA design in beamforming design. In addition, the analog precoding design with a fully-connected structure is always superior to that with a sub-connected structure, since all RF chains in the fully-connected structure are connected to all antenna elements through PS. It is noteworthy that when the power is less than -5 dBm, the WSR value of the RA scheme is better than that of the fully-digital beamforming design. This is due to the consideration of antenna aperture loss; at low transmit power, the improvement of system performance mainly relies on antenna rotation. However, as the transmit signal power increases, the gain from beamforming becomes increasingly prominent and becomes the dominant factor driving the improvement of system performance, thus making the WSR lower than that of the fully-digital beamforming scheme.
@@
Fig. 6 shows the relationship between the sensing rate and the maximum transmit power $P_{max}$ under different schemes. We set $\omega_{s}=1$, whic means the ISAC system is sensing-focused. As can be seen from the figure, as $P_{max}$ increases, the sensing rate of each schemes shows an upward trend. It can be intuitively observed from the figure that array rotation brings significant gain to sensing performance, and at the same time, the impact of different precoding structures on the sensing rate can also be clearly seen. Among them, the sensing rate of the fully digital beamforming scheme is lower than that of the RA scheme. This is because in the scenario where only a single target is sensed, the gain generated by array rotation is significantly greater than the beamforming gain.
Fig. 6 illustrates the the sum of communication rate versus the maximum transmit power $P_{max}$ at a communication-focused ISAC with $\omega_{c}=1$. A similar trend with Fig. 6 can be observed, and the performance improvement on communication is more significant than that of sensing. Notably, when $P_{max}>5$dbm, the performance of the RA scheme with a sub-connected beamforming is lower than that of the FPA scheme with fully-connected beamforming. Because in the multi-user scenario, as the transmit power increases, the antenna rotation gain in the sub-connected structure is smaller than the beamforming gain in the fully-connected structure.
@@
Figure 6: Sum of communication rate versus $P_{max}$. $N_{t}=N_{r}=16$, $M=4$, $\varpi_{c}=1$.
@@
Figure 8: WSR versus $M$. $N_{t}=N_{r}=64$, $D=10\lambda$, $\varpi_{s}=0.7$.
In Fig. 8, we analyze the trade-off between sensing an communication. We set $P_{max}=0$ dbm, $D=10\lambda$. By adjusting the sensing rate weight $\omega_{s}$ and the communication weight $\omega_{c}$, the trade-off curves between the sensing rate and the communication rate are ploted. To analyze the impact of RA and sub-connected structure $M$ on system performance, we consider three typical ISAC system configurations: 1) Number of transmit/receive antennas $N_{t}=N_{r}=8,M=2$; 2) $N_{t}=N_{r}=8,M=4$; 3) $N_{t}=N_{r}=16,M=2$. As shown in Fig. 8, under the condition of the same number of antennas and $M$, the ISAC system adopting RA outperforms the FPA system in terms of the performance trade-off between communication and sensing. This is because RA can more accurately focus the beam on the desired position. Specifically, with a fixed number of antennas, more RF chains can enhance the DoF of beamforming but will increase hardware costs. While with a fixed M, more antennas need more RF chains to achieve an improvement in overall system performance. It is worth noting that when the sensing weight $\omega_{s}$ is large, the performance of RA-aided ISAC system schemes with fewer antennas is better than that of FPA schemes with more antennas. The reason is that RA can significantly improve the sensing performance of the system. As the $\omega_{s}$ decreases, the system is communication-focused. At this point, the performance of RA schemes with fewer antennas gradually becomes inferior to that of FPA scheme with more antennas. This is bacause the combination of an increased number of antennas and more RF chains provides higher DoF for beamforming, which is more suitable for the requirements of communication-focused ISAC. Hence, appropriate weighted factors and configuration parameters need to be considered in partial engineering.
Finally, Figure. 8 illustrates the relationship between the WSR and the number of RF chains. We consider three system configurations: fully-connected hybrid beamforming, sub-connected hybrid beamforming, and fully digital beamforming. We set $N_{t}=N_{r}=64,A=10\lambda,\omega_{c}=0.7$. It can be observed that as the $M$ increases, the number of RF chains decreases correspondingly, and the WSR under the hybrid beamforming design show a downward trend. This can be explained from two perspectives. On one hand, an increased number of RF chains provides more DoFs for the downlink beamfocusing and thus mitigates the inter-user interference among CUs; On the other hand, with an increased number of RF chains, more sophisticated probing signal design can be achieve to realize better positioning performance. When the number of RF chains drops to a specific number, the performance of fully-connected and sub-connected hybrid precoding tends to be close, and at this time, RA becomes the main contributing factor to the system performance improvement.
@@ V Conclusion
In this paper, we study an RA-aided sub-connected hybrid beamforming design for ISAC systems. By jointly optimizing the angle of the rotatable array and the hybrid beamforming vector, we maximize the system’s achievable WSR for both communications and sensing. We proposed a alternating optimization based framework for solving this non-convex optimization problem. We transform the objective function using the FP method, and alternately solve five sub-problems until the AO algorithm converges. We adopt the MMSE filtering method to optimize the receive beamformer. Then we derive the closed-form expressions for updating the digital and analog beamforming matrices. We propose a GA-based method to optimize the angle of the rotatable array. Simulation results verified the effectiveness of the proposed algorithm and the advantages of the considered RA-aided sub-connected structure ISAC scheme compared to FPA-based schemes. Remarkably, under some weight parameters, the performance of RA-aided ISAC system equipped with 8 transmit/receive antennas and 4 RF chains is outperformed to that of FPA ISAC system equipped with 16 transmit/receive antennas and 8 RF chains. This finding provides valuable insights for reducing hardware costs in engineering applications. In addition, apart from the sensing rate, this paper also derives and analyzes the sensing beampattern. The results show that RA can effectively optimize the beamfocusing effect and improve the sensing accuracy.

### v2

@@ III-F Overall Algorithm
The total complexity of Algorithm 2 stems from the solving the variables $\{\mathbf{u},\mathbf{W},\mathbf{F},\mathbf{\Lambda},\mathbf{\mu},\bm{\xi}^{c},\bm{\xi}_{s}\}$. The computational complexity for calculating receive beamformer $\mathbf{u}$ is $\mathcal{O}(N_{r}^{3})$ due to the matrix inversion in (24). Similarly, the computational complexities for optimizing transmit digital beamformer $\mathbf{W}$ and transmit analog beamformer $\mathbf{F}$ are both $\mathcal{O}(N_{t}^{3})$. To update the auxiliary variables $\mathbf{\mu},\bm{\xi}^{c}$ and $\bm{\xi}_{s}$, the complexities are $\mathcal{O}(KN_{t})$, $\mathcal{O}(KN_{t}+K^{2})$ and $\mathcal{O}(N_{t}K^{2})$, respectively. The complexity for updating $\mathbf{\Lambda}$ mainly stems from the derivative calculations (51) for the $L$ elements in $\mathbf{\Lambda}$. Thus, the computational complexity of $\mathbf{\Lambda}$ is $\mathbf{O}(I_{GA}LN_{t}^{2}K^{2})$, where $I_{GA}$ is the number of iterations for gradient ascent. Considering the dominant computational steps, the total computational complexity of Algorithm 2 is $\mathcal{O}(I_{AO}I_{GA}LN_{t}^{2}K^{2})$, where $I_{AO}$ denotes the number of iterations of the AO algorithm.
@@ IV Numerical results
In this section, we elaborate on the simulation setup and discussion of the numerical results to evaluate the performance of the proposed algorithm.
@@ IV-A Simulation Setup
In the simulation, the BS is the origin of the global coordinate system. The center coordinates of the transmitting uniform planar array (UPA) and the receiving UPA are set to $0,0,-d_{0}]^{T}$ and $0,0,d_{0}]^{T}$, respectively. The BS is equipped with $N_{t}$ transmit antennas, $N_{r}$ receive antennas, and $N_{RF}$ RF chains. Each RF chain is connected to $M$ antennas via analog phase shifters. The antenna’s positions are set according to the transmit and receive UPAs with the largest achievable apertures $D\times D$. For each UPA, the panel size is $D\times D$ with $D=10\lambda$. The BS operates at a carrier frequency of 30GHz. It serves $K$ = 4 users and senses $s=1$ target, with the interference of C=C=3 clutters. The users and clutters are randomly distributed in front of the BS, i.e., $\phi_{d}\sim\mathcal{U}(-\frac{\pi}{2},\frac{\pi}{2})$, and $\theta_{d}\sim\mathcal{U}(0,\pi)$ for d $\in\{1,\dots,K+C\}$. The sensing target is assumed to be located at $(\frac{\pi}{3},\frac{\pi}{3})$. The reflection coefficients of both the target and the clutters, $\zeta_{k}$ and $\zeta_{s}$ , are set to 2 dB. In addition, the power budget at the BS is set as $P_{max}$ = 0 dBm. The average noise power is $\sigma^{2}_{k}=\sigma^{2}_{s}=$ -80 dBm. The penalty factor $\eta=$ 10. The preliminary step size $\kappa=100$ and the minimum step size $\kappa_{min}$ = 0.01. It is worth noting that, unless otherwise specified, the optimized rotation variable $\mathbf{\Lambda}$ includes all rotation parameters of the TP and RP, i.e., $\mathbf{\Lambda}=[\alpha^{\mathrm{t}},\beta^{\mathrm{t}},\gamma^{\mathrm{t}},\alpha^{\mathrm{r}},\beta^{\mathrm{r}},\gamma^{\mathrm{r}}]$. We perform 200 Monte Carlo simulations for the following results. The simulation parameters for the proposed system are listed in Table. I
@@
To verify the effectiveness of the proposed scheme with Algorithm 2, we introduce six different schemes as follows,
  1. 1.
    The TP and RP have fixed rotation angles, with RF chains adopting a sub-connected structure (FPA, SC).
  2. 2.
    The TP and RP have fixed rotation angles, with RF chains adopting a fully-connected structure (FPA, FC).
  3. 3.
    The TP and RP have fixed rotation angles, using a fully-digital beamformer (FPA, FD).
  4. 4.
    The TP can execute 3D rotations and RP has fixed rotation angles, with RF chains adopting a sub-connected structure (Transmit RA & receive FPA, SC).
  5. 5.
    The TP has fixed rotation angles and RP can execute 3D rotations, with RF chains adopting a sub-connected structure (Transmit FPA & receive RA, SC).
  6. 6.
    Both the TP and RP can execute 3D rotations, with RF chains adopting a sub-connected structure (RA, SC).
@@ IV-B Convergence Evaluation
First, we evaluate the convergence of the proposed AO algorithm under different numbers of RF chains in Fig. 2. As can be seen, the algorithm converges rapidly in all scenarios, and the objective value tends to stabilize after 20 iterations. Moreover, the WSR increases as the number of RF chains increases. This is because the performance of the analog beamformer $\mathbf{F}$ is constrained by the number of RF chains. Specifically, the hybrid beamformer can achieve a higher DoFs to further enhance the overall system performance when the number of RF chains increases.
@@ IV-C Beamfocusing of Algorithm 2
To verify the effect of Algorithm 2 on improving sensing performance, the optimized beampattern by Algorithm 2 is compared with the optimized beampattern by FPA scheme. It intuitively demonstrates the advantages of RA in the sensing function of ISAC system proposed in this paper. The beampattern at the angle ($\phi,\theta$) is computed as
@@
Set $\omega_{s}=1$ to eliminate the interference from CUs. The number of transmit antennas and receive antennas are both $N_{t}=N_{r}=16$, the number of RF chains is $N_{RF}=4$, and the target is assumed to be located at ($60^{\circ},60^{\circ}$). Fig.3(a) shows the optimized beampattern under the FPA scheme, which indicates that the beam fails to focus well on the target position. In Fig.3(b), the side length of the UPA is set to $D=5\lambda$, and it is observed that under the RA-aided scheme, the beam can focus precisely on the desired target position with extremely lower interference to other positions. In the far-field scenario, as the side length of the UPA increases as shown in Fig.3(c), the width of the beam main lobe narrows gradually. However, due to the increase in the spacing between antennas in the array, the array periodicity is enhanced and the number of grating lobes increases, resulting in more interference.
@@
Figure 3: Beampattern under different configurations.
@@ IV-D Simulation Results
Fig. 4 compares the relationship between the WSR and the maximum transmit signal power $P_{max}$ among different schemes. With the increase of transmit signal power, the WSR values of both the RA and FPA schemes keep rising, which is because a higher transmit power can enhance the signal strength. It is noteworthy that the performance of the RASC scheme is superior to that of all other schemes. Specifically, when the transmit power is 0 dBm, it achieves performance improvements of $16.96\%$, $60.42\%$, and $82.39\%$ compared with the FPAFD, FPAFC and FPASC schemes, respectively. This is attributed to the RA scheme’s ability to achieve effective improvement in channel conditions and exhibit higher flexibility in beamforming design. When only the RP is rotatable, the system can only optimize the echo channel but cannot improve the DoFs of beamforming at transmitter, resulting in limited performance improvement. When only the TP is rotatable, although it can effectively enhance user performance compared with the FPASC scheme, it is still inferior to the proposed RASC scheme. The latter realizes the optimization of the echo channel and the improvement of beamforming capability simultaneously through the rotation of TP and RP, which fully proves the effectiveness of the proposed scheme in improving the performance of the ISAC system.
@@
Fig. 6 shows the relationship between the sensing rate and the maximum transmit power $P_{max}$ under different schemes. We set $\omega_{s}=0.8$, which means the ISAC system is sensing-focused. As $P_{max}$ increases, the sensing rate of each scheme shows an upward trend. It can be intuitively observed from the figure that array rotation brings significant gain to sensing performance, and at the same time, the impact of different precoding structures on the sensing rate can also be clearly seen. In particular, although the system is sensing-focused, the sensing performance of the RARoT scheme is superior to that of the RARoR scheme, due to the rotation of the TP can effectively enhance the beamforming capability.
Fig. 6 illustrates the sum of communication rate versus the maximum transmit power $P_{max}$ at a communication-focused ISAC with $\omega_{c}=0.8$. A similar trend with Fig. 6 can be observed, and the performance improvement on communication is more significant than that of sensing. Notably, when the system is communication-focused, compared with FPASC scheme, the scheme that only rotates the RP will instead lead to a decrease in the system’s communication capability. This is because although the rotation of the RP can effectively improve the system’s sensing capability, it will cause the system to allocate more transmit power to sensing during the resource trade-off process, thereby resulting in a slight decline in the communication performance.
@@
Figure 6: Sum of communication rate versus $P_{max}$. $N_{t}=N_{r}=16$, $M=4$, $\varpi_{c}=0.8$.
@@
Figure 8: WSR versus $M$. $N_{t}=N_{r}=64$, $D=10\lambda$, $\varpi_{c}=0.5$.
Fig. 8 analyzes the trade-off between sensing and communication with $P_{max}=0$ dBm and $D=10\lambda$. By adjusting the sensing rate weight $\omega_{s}$ and the communication rate weight $\omega_{c}$, we obtain the trade-off curves between the sensing rate and the communication rate. To analyze the impact of RA and sub-connected structure on system performance, let $N_{t}=N_{r}=16$. Meanwhile, two sub-connection structures ($N_{RF}=4,N_{RF}=8$) and a fully-connected structure $(N_{RF}=8)$ are considered. As shown in Fig. 8, RA can significantly improve system performance compared with FPA. Under the FPA scheme, the system performance of the precoding structure is inferior to that of the fully-digital beamformer. But the performance advantage of the fully-digital beamformer is achieved at the cost of higher hardware expenses. For the precoding structure, the more RF chains, the better the system performance. When the number of RF chains is the same, the system performance of the fully-connected structure is superior to that of the sub-connected structure. However, the fully-connected structure uses more PSs, which makes the system hardware more complex. It is worth noting that under the RA scheme, the system performance is similar between the cases with 4 RF chains and 8 RF chains when the communication weight $\omega_{c}=0$. This indicates that the number of RF chains has almost no impact on the target SCNR. However, $\omega_{c}$ increases and the system is gradually communication-focused, the performance of the system with 8 RF chains is significantly superior to that with 4 RF chains. Hence, appropriate weighted factors and configuration parameters need to be considered in partial engineering.
Finally, Figure. 8 illustrates the relationship between the WSR and the number of RF chains. The number of 8 $\times$ 8 UPAs is $N_{t}=N_{r}=64$. It can be observed that as the $N_{RF}$ increases, the WSR under the hybrid beamforming design shows an upward trend. This can be explained from two perspectives. On one hand, an increased number of RF chains provides more DoFs for the downlink beamfocusing and thus mitigates the inter-user interference among CUs; On the other hand, with an increased number of RF chains, a more sophisticated probing signal design can be achieved to realize better positioning performance. When the number of RF chains exceeds 8, the system performance of the RASC scheme is superior to that of the FPA fully-digital beamforming scheme. This result indicates that the proposed RA scheme can significantly reduce the number of RF chains while still achieving performance that surpasses that of the fully-digital beamforming scheme.
@@ V Conclusion
In this paper, we study an RA-aided sub-connected hybrid beamforming design for ISAC systems. By jointly optimizing the angle of the rotatable array and the hybrid beamforming vector, we maximize the system’s achievable WSR for both communications and sensing. An AO-based framework is proposed to solve this non-convex optimization problem. We transform the objective function using the FP method, and alternately solve five sub-problems until the AO algorithm converges. Firstly, the MMSE filtering method is adopted to optimize the receive beamformer. Moreover, the closed-form expressions is derived to update the digital and analog beamforming matrices. Then we derived the derivative expression of the objective function with respect to the rotation angle through research and analysis. Based on this expression, a GA method is employed to optimize the angle parameters of the rotatable array. Simulation results verified the effectiveness of the proposed algorithm and the advantages of the considered RA-aided sub-connected structure ISAC scheme compared to FPA-based schemes. Remarkably, under some weight parameters, the performance of RA-aided hybrid beamforming ISAC system equipped with 64 transmit/receive antennas and 8 RF chains is outperformed to that of FPA fully-digital ISAC system equipped with 64 transmit/receive antennas. This finding provides valuable insights for reducing hardware costs in engineering applications. In addition, apart from the sensing rate, this paper also derives and analyzes the sensing beampattern. The results show that RA can effectively optimize the beamfocusing effect and improve the sensing accuracy.